OpenStudy (anonymous):

Answer needed fast! please? 5 miles is approximately equal to 8 kilometres. Roughly how many kilometres are equal to 35 miles? How many miles are equal to 112 km?

OpenStudy (anonymous):

I need to show how i got my answer.

OpenStudy (anonymous):

56 to 35

OpenStudy (anonymous):

5 miles-8 kilometres 1 mile - (8/5) km 35 miles - (8/5)x 35 = 56 km 112km - (5/8) x 112 = 70 miles

OpenStudy (anonymous):

or use a proportion 5/8 = 35/ x = 56

OpenStudy (anonymous):

and 5/8 = x/112

OpenStudy (anonymous):

Asgar.. Why is it's 112km - *(5/8)*? x 112 = 70 miles why not (8/5)?

OpenStudy (anonymous):

what you do is set up the first proportion

OpenStudy (anonymous):

you were given 5 miles is equal to 8 kilometres

OpenStudy (anonymous):

now 35 is on top next to the 5 because 35 is in miles and you are looking for kilometres

OpenStudy (anonymous):

so it is 56

OpenStudy (anonymous):

Hmmmmmmmmmmmmmmmmmmm! Got it! Thanks Guys!

OpenStudy (anonymous):

you understand now you need any other help?